OpenStudy (anonymous):

What is the area of an equilateral triangle with sides of 6 inches.?

OpenStudy (anonymous):

A = sq. rt3 x a^2/ 4 = 1.732 x 6^2/ 4 = 1.732 x 36 /4 = 15.588

OpenStudy (anonymous):

oh my goodness, thank you. xD I got that a couple times, but every time I did it I would get a different answer.